1) The U.S. GDP in 2008 was $14,291.5 billion, and the previous year the GDP was $14, 028.7 billion. What was the rate of growth in GDP in 2008? O $283 billion 0 1.84% О 7.02% О 1.87% $14, 160.1 billion2) Use the table to answer the questions. Define productivity as GDP per worker, and please enter all answers as numerical values rounded to 3 decimal places, and not as percentages (e.g., 0.103 instead of 10.3%). GDP Population Workers 121 241 Year 1071 175 Year 2 1183 279 What is the GDP growth rate? What is GDP per capita growth rate? What is the productivity growth rate?
Answer #1

Answer : 1) The answer is option D.

Growth rate of year 2008 = [(GDP of year 2008 - GDP of previous year) / GDP of previous year] × 100

=> Growth rate of year 2008 = [(14291.5 - 14028.7) / 14028.7] × 100 = 1.87%

Therefore, the growth rate of year 2008 is 1.87%.

2) GDP per worker productivity : GDP per worker productivity is the productivity level of per worker which is included in GDP. GDP per worker productivity is calculated by the following formula :

GDP per worker productivity = GDP / Worker .

a) GDP growth rate = (GDP of year 2 - GDP of year 1) / GDP of year 1

=> GDP growth rate = (1183 - 1071) / 1071 = 0.105

Therefore, the GDP growth rate is 0.105 .

b) GDP per capita for year 1 = GDP / Population = 1071 / 175 = 6.120

GDP per capita for year 2 = 1183 / 279 = 4.240

GDP per capita growth rate = (GDP per capita of year 2 - GDP per capita for year 1) / GDP per capita for year 1

=> GDP per capita growth rate = (4.240 - 6.120) / 6.120 = - 0.307

Therefore, GDP per capita growth rate is - 0.307 .

c) GDP per worker productivity for year 1 = GDP / Worker = 1071 / 121 = 8.851

GDP per worker productivity for year 2 = 1183 / 241 = 4.909

Productivity growth rate = (GDP per worker productivity of year 2 - GDP per worker productivity of year 1) / GDP per worker productivity of year 1 = (4.909 - 8.851) / 8.851 = - 0.445

Therefore, the productivity growth rate is - 0.445 .